The diameters of the bases of two cones are epual.the ratio of their slant heights is 5:4 if curved surface of the smaller cone is 400cm2 then find the curved surface of the bigger cone.
Answer by dkppathak(439)

The diameters of the bases of two cones are epual.the ratio of their slant heights is 5:4 if curved surface of the smaller cone is 400cm2 then find the curved surface of the bigger cone.
Solution let Curved surface area of cone Is X cm^2
Let radius of cone are r1 and r2
r1:r2=1:1 (given)
let slant height are L1 and L2
L1:L2=5:4 (1)
curved surface area of cone =pie rL
ratio of curves surface are of cone = pie r L1/pie r L2
ratio of curved surface area =L1/L2
L1:L2=X:400 from (1)
5/4=X/400
4X= 2000
X=2000/4
X=500 Cm^2
curves surface area of cone 500 cm^2